# Future collider constraints on top-quark operators

### v2

@@ 1 Introduction
Precision measurements of differential cross sections and asymmetries Perello Rosello and Vos (2016) in top-quark pair production provide bounds on the top-gluon vertex and $q\bar{q}t\bar{t}$ operators. Measurements of single top-quark production and top-quark decay constrain the $tWb$ vertex. Run 2 of the LHC opened up rare associated production processes that yield direct bounds on the top-quark couplings to neutral bosons, i.e. the $t\bar{t}\gamma$, $t\bar{t}Z$ and $t\bar{t}H$ vertices. The observation of four-top-quark production Aad et al. (2021, 2023a); Hayrapetyan et al. (2023) and studies of $t\bar{t}b\bar{b}$ Sirunyan et al. (2020) even yield the first direct constraints on four-heavy-quark operators. The least constrained set of operators is formed by two-lepton-two-heavy-quark ($l^{+}l^{-}Q\bar{Q}$) operators. There are five independent $l^{+}l^{-}Q\bar{Q}$ operators involving top quarks (15 if lepton flavour universality is dropped) that are only very loosely constrained by measurements of $pp\rightarrow t\bar{t}l^{+}l^{-}$ production collaboration (2023).
@@
In this paper, we update and complete the studies reported in Refs. Durieux et al. (2018, 2019, 2022); de Blas et al. (2022a); Agashe et al. (2022), updating and extending these results. We study new LHC measurements of the entanglement of top-quark pairs The CMS collaboration (2024a, b) and include projections for $t\bar{t}l^{+}l^{-}$ for the HL-LHC. We adopt up-to-date scenarios for all future electron-positron collider projects, and include a study of the potential of a muon collider.
@@ 2 SMEFT basis
The decoupling theorem Appelquist and Carazzone (1975) guarantees that the contributions of the heavy degrees of freedom to the physical amplitudes are suppressed by the inverse of their masses up to logarithmic corrections. This allows us to condensate the NP effects into higher dimensional operators built with the SM particle content. In the SMEFT the NP is assumed to be invariant under the same gauge symmetries as the SM Lagrangian before the electroweak symmetry breaking. Imposing also that the accidental symmetries of the SM are conserved (baryon and lepton number) the SMEFT Lagrangian can be written as
@@
were the operators $O_{i}$ are built from SM fields and the coefficients $C_{i}$ are known as the Wilson coefficients (WC).
@@
The linear terms arise from the interference of the NP contributions with the SM and the quadratic terms come from squaring the NP contributions. In our work we will only include the leading-order (LO) contribution from the dimension-six operators. The quadratic terms are of order $\Lambda^{-4}$, the same order as the interference of the dimension-eight operators with the SM. A complete treatment of the $\mathcal{O}(\Lambda^{-4})$ terms is currently not possible. Several groups have developed recipes to truncate the expansion and define uncertainties to ensure validity of the SMEFT Brivio et al. (2022), but there is no consensus on which approach to adopt. In our analysis, we include only the linear terms proportional to $\Lambda^{-2}$. In analyses of the LHC and HL-LHC projections, quadratic terms can be sizable, and our results are therefore conservative in comparison with other groups. Quadratic terms generally play a minor role at lepton colliders at energies up to several TeV . Therefore, our projections for the Higgs factory would not be altered significantly by the inclusion of $\Lambda^{-4}$ terms.
@@ 3 Current Constraints
In Fig. 1 and Tab. 4 we show the 95% probability limits for the top-quark operators using the current LHC data. All Wilson coefficients are found to be compatible with the SM value of 0 at 95% probability, reflecting the good agreement of the measurements of Tab. 3 with the SM predictions. Regarding the goodness-of-fit, we find that the SM has a $\chi^{2}$ per degree of freedom of 0.76, whereas the SMEFT fit yields a significantly improved value of 0.56.111When determining the number of degrees of freedom we have considered each bin as an observable even though this value may be reduced due to correlations among the bins.
The tightest bounds are obtained on the coefficients of two-fermion operators, where 95% probability limits range from 0.25 TeV -2 for $C_{tW}/\Lambda^{2}$ to 5 TeV -2 for $C_{t\varphi}$. These bounds are driven by precise differential cross section measurements in top-quark decay, single top-quark production and associated production processes $t\bar{t}H$, $t\bar{t}Z$ and $t\bar{t}\gamma$. In particular, the studies of rare associated top-quark production processes have evolved from observation to precision measurements. It is illustrative to contrast our result with the first fits of these coefficients at the start of Run 2 in Ref. Buckley et al. (2016), where the data still provided insufficient constraints for a global analysis.
@@
The fit does not yield meaningful bounds on two-quark two-lepton operators, as the measurements of Tab. 3 have insufficient sensitivity for the four degrees of freedom, and therefore these WC have not been considered for this fit. Indeed, even when setting the NP scale as low as 1 TeV the constraints of these operators in the global fit are too weak to lay within the perturbative regime. Dedicated analyses of $t\bar{t}l^{+}l^{-}$ at the LHC with $m_{ll}$ above the $Z$-boson mass window may provide additional constraints on these operators. The analysis of Ref. collaboration (2023) finds bounds of ${\mathcal{O}}(1~{}{\mathrm{Te\kern-1.00006ptV}}\,^{-2})$. These bounds, however, strongly rely on terms proportional to $\Lambda^{-4}$. The Run 2 data set cannot constrain these operator coefficients in a global analysis based on a linear parametrisation.
@@ 4.3 Associated production with additional charged leptons
The analysis considers several degrees of freedom that correspond to operators with two charged leptons and two top quarks, listed in the lowermost box of Tab. 2. The LHC experiments can constrain their WC through the study of top-quark production in association with additional charged leptons. Bounds of order one are reported from a SMEFT fit to LHC Run 2 data by the CMS experiment collaboration (2023).
The projection of HL-LHC bounds on these operator coefficients is based on an extrapolation of the measurements in $pp\rightarrow t\bar{t}Z$ production The ATLAS collaboration (2021). The analysis targets events with three or four isolated charged leptons (electrons and muons). Instead of the events with $m_{ll}\sim m_{Z}$ used in Ref. The ATLAS collaboration (2021), here we focus on off-shell events, where the invariant mass of the di-lepton system lies outside the $Z$-boson mass window. Events are binned in four $m_{ll}$ bins: [100-120] GeV , [120-140] GeV , [140-180] GeV and $m_{ll}>$ 180 GeV  A. Gutierrez Camacho, M. Moreno Llácer and M. Vos (2022). The observed event yields of the Run 2 analysis range from $\sim$ 30 events in the lowest bin to about 10 events for the highest bin. Individual bounds based on Run 2 data with quadratic parametrisations of $\Lambda^{-2}$ and $\Lambda^{-4}$ terms yield results similar to Ref. collaboration (2023), where the sensitivity is driven by the $\Lambda^{-4}$ terms and most pronounced in the highest mass bin.
@@ 4.4 Projections
Marginalised bounds remain significantly less precise than individual bounds, even after the complete HL-LHC program, due to “blind directions”, i.e. unresolved correlations between the large number of degrees of freedom. This is confirmed in fits to the top-quark sector by other groups Brivio et al. (2020); Hartland et al. (2019) and even in global Higgs/EW/top fits Ethier et al. (2021); Ellis et al. (2021). The blind directions have a less pronounced effect in quadratic fits, where dimension-six-squared terms proportional to $\Lambda^{-4}$ are included Ethier et al. (2021), as strictly positive contributions cannot cancel each other.
@@ 5 Prospects for $e^{+}e^{-}$ colliders
Machine P($e^{+}$, $e^{-}$) Energy Luminosity Reference
ILC (±$(\pm 30\%,\,\mp 80\%)$ 250 GeV 2 ab-1 A. Aryshev et al. () (editors)
@@ 3 TeV 5 ab-1
FCC-$ee$ Unpolarised Z-pole 150 ab-1 Bernardi et al.
240 GeV 5 ab-1
350 GeV 0.41 ab-1
365 GeV 2.65 ab-1
@@ 5.4 Projections
To assess the potential of electron-positron colliders, the fit is repeated by adding the projected measurements in $e^{+}e^{-}\rightarrow b\bar{b}$, $e^{+}e^{-}\rightarrow t\bar{t}$ and $e^{+}e^{-}\rightarrow t\bar{t}H$ production. These processes are accessible in successive energy stages of the electron-positron collider. The impact of each process and energy stage is made clear for the example of the ILC in Fig. 7.
@@
The projections for the top-quark Yukawa, translating $\delta y_{t}=-\frac{v^{2}}{\Lambda^{2}}C_{t\varphi}$, are also presented in Tab. 6. Numbers for lepton colliders are based on an extrapolation in Ref. de Blas et al. (2022a) of detailed studies in Refs. Price et al. (2015); Abramowicz et al. (2019). We emphasise the effect of increasing the ILC from 500 GeV to 550 GeV at the same integrated luminosity of 4 ab-1. The increase in centre-of-mass energy translates into a reduction of the uncertainty by a factor 1.5 for the individual limit thanks to the increase of a factor of 3 in the cross section $e^{+}e^{-}\rightarrow t\bar{t}H$. Higher-energy operation can improve the limit further, as demonstrated by the ILC1000 scenario, with 8 ab-1.
@@
$\delta y_{t}$ Global 16% 5.6% 3.4% 2.8% 1.4% 3.0%
Indiv. 13% 4.1% 2.9% 2.4% 1.4% 2.7%
Table 6: Uncertainties for the top-quark Yukawa coupling at 68% probability for different scenarios, in percent. The ILC500, ILC550 and CLIC scenarios also include the HL-LHC. The ILC1000 scenario includes also ILC500 and HL-LHC. Numbers for lepton colliders are based on an extrapolation in Ref. de Blas et al. (2022a) of detailed studies in Refs. Price et al. (2015); Abramowicz et al. (2019).
@@ 6 Lepton colliders in the 3–30 TeV  range
In this section, we study the impact on the top-quark sector of the SMEFT of lepton colliders operated at very high energy, with $\sqrt{s}$ in the range from several to several tens of TeV. This regime could be unlocked by a muon collider, that presents unique benefits (and technological challenges) in comparison with classical electron-positron machines. The Snowmass process provides several compelling reasons to pursue the construction of a Muon Collider Narain et al. (2022), and the P5 Asai et al. (2023) report has endorsed R&\&&D efforts. Unlike circular electron-positron colliders, muon colliders can reach very high centre-of-mass energies without significant energy losses due to synchrotron radiation.
@@ 7 Conclusion
We also provide projections for the high-luminosity phase of the Large Hadron Collider and future colliders. HL-LHC projections are obtained adopting the “S2” scenario that is also used for Higgs measurements Cepeda et al. (2019). In this scenario, experimental uncertainties are strongly reduced $(\propto 1/\sqrt{L_{\text{int}}})$, while theory and modelling uncertainties see a more modest improvement (by a factor 1/2). The bounds improve by a factor 2-3 for most of the operator coefficients. In the projections, we extend previous studies Durieux et al. (2019) by considering also seven operators with two charged leptons and two top quarks, enlarging the total number of Wilson coefficients to 29. In principle, these can be constrained by measurements of $t\bar{t}l^{+}l^{-}$ production with $m_{ll}>m_{Z}$. However, we find that no meaningful global bounds are obtained with a parametrisation based on terms proportional to $\Lambda^{-2}$
A lepton collider operated above the ${t\bar{t}}$ production threshold has excellent potential to constrain the coefficients of bottom- and top-quark operators. The projections used in this paper are based on full-simulation study of optimal observables in Ref. Durieux et al. (2018), with up-to-date operating scenarios for electron-positron Higgs/top/EW factories. Individual bounds on the coefficients $C_{tZ}$ and $C_{tW}$ of the dipole operators reach and on $C_{\varphi Q}^{(1)}$ and $C_{\varphi Q}^{(3)}$ reach 10-2 TeV -2 with a single run with an integrated luminosity of 4 ab-1 at $\sqrt{s}=$ 500 GeV . The bounds on the two-lepton-two-heavy-quark operators even reach 10-3 TeV -2. The global bounds are limited by degeneracies between the two-fermion and four-fermion operators, that can be lifted to a good extent by including measurements at two different centre-of-mass energies.
@@ Acknowledgments
The authors acknowledge the work of Master students Abel Gutierrez Camacho, Belén Durán González and Pablo Copete Garrido in the development of this analysis. F.C.G is supported by the Presidential Society of STEM Postdoctoral Fellowship at Case Western Reserve University and by the Ministerio de Ciencia, Innovación y Universidades, Spain, through a Beatriz Galindo Junior grant BG23/00061. The work of VM has been supported by the Italian Ministry of Research (MUR) under the grant PRIN20172LNEEZ, by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 949451) and a Royal Society University Research Fellowship through grant URF/R1/201553. The work of MML is supported by the Science and Technology Facilities Council [grant number ST/X005941/1]. MMLL acknowledges the support received from the Spanish Ramón y Cajal programme (RYC2019-028510-I). Her work is also supported by the projects ASFAE/2022/010 and CIPROM/2022/70 (Generalitat Valenciana), and PID2021-124912NB-I00 (Spanish Ministry MICIN). The work of MV is supported by the Spanish ministry for science under grant number PID2021-122134NB-C21, by the Generalitat Valenciana under PROMETEO grant CIPROM/2021-073, and by CSIC under grant ILINKB20065. The group members of IFIC in Valencia received support from the Severo Ochoa excellence programme.

### v3

@@ 1 Introduction
Precision measurements of differential cross sections and asymmetries Perello Rosello and Vos (2016) in top-quark pair production provide bounds on the top-gluon vertex and $q\bar{q}t\bar{t}$ operators. Measurements of single top-quark production and top-quark decay constrain the $tWb$ vertex. Run 2 of the LHC opened up rare associated production processes that yield direct bounds on the top-quark couplings to neutral bosons, i.e. the $\bar{t}\gamma, tt$$\bar{t}Z$ and $t\bar{t}H$ vertices. The observation of four-top-quark production Aad et al. (2021, 2023a); Hayrapetyan et al. (2023) and studies of $t\bar{t}b\bar{b}$ Sirunyan et al. (2020) even yield the first direct constraints on four-heavy-quark operators. The least constrained set of operators is formed by two-lepton-two-heavy-quark ($l^{+}l^{-}Q\bar{Q}$) operators. There are five independent $l^{+}l^{-}Q\bar{Q}$ operators involving top quarks (15 if lepton flavour universality is dropped) that are only very loosely constrained by measurements of $pp\rightarrow t\bar{t}l^{+}l^{-}$ production The CMS collaboration (2023); Aad et al. (2025).
@@
In this paper, we update and complete the studies reported in Refs. Durieux et al. (2018, 2019, 2022); de Blas et al. (2022a); Agashe et al. (2022). We study new LHC measurements of the entanglement of top-quark pairs The CMS collaboration (2024a, b) and include projections for $t\bar{t}l^{+}l^{-}$ for the HL-LHC. We adopt up-to-date scenarios for all future electron-positron collider projects, and include a study of the potential of a muon collider.
@@ 2 SMEFT basis
The decoupling theorem Appelquist and Carazzone (1975) guarantees that the contributions of the heavy degrees of freedom to the physical amplitudes are suppressed by the inverse of their masses up to logarithmic corrections. [^1] Strictly speaking, this does not generally hold in spontaneously broken gauge theories when the masses of the heavy particles arise from the symmetry breaking itself. In such cases, non-decoupling effects may occur. This allows us to condensate the NP effects into higher dimensional operators built with the SM particle content. In the SMEFT the NP is assumed to be invariant under the same gauge symmetries as the SM Lagrangian before the electroweak symmetry breaking. Imposing also that the accidental symmetries of the SM are conserved (baryon and lepton number) the SMEFT Lagrangian can be written as
@@
where the operators $O_{i}$ are built from SM fields and the coefficients $C_{i}$ are known as the Wilson coefficients (WC).
@@
The linear terms arise from the interference of the NP contributions with the SM and the quadratic terms come from squaring the NP contributions. In our work we will only include the leading-order (LO) contribution from the dimension-six operators. The quadratic terms are of order $\Lambda^{-4}$, the same order as the interference of the dimension-eight operators with the SM. A complete treatment of the $\mathcal{O}(\Lambda^{-4})$ terms is currently not possible. Several groups have developed recipes to truncate the expansion and define uncertainties to ensure validity of the SMEFT Brivio et al. (2022), but there is no consensus on which approach to adopt. In our analysis, we include only the linear terms proportional to $\Lambda^{-2}$. In analyses of the LHC and HL-LHC projections, quadratic terms can be sizeable, and our results are therefore conservative in comparison with other groups. Quadratic terms generally play a minor role at lepton colliders at energies up to several TeV . Therefore, our projections for the Higgs factory would not be altered significantly by the inclusion of $\Lambda^{-4}$ terms.
@@ 3 Current Constraints
In Fig. 1 and Tab. 4 we show the 95% probability limits for the top-quark operators using the current LHC data. All Wilson coefficients are found to be compatible with the SM value of 0 at 95% probability, reflecting the good agreement of the measurements of Tab. 3 with the SM predictions. Regarding the goodness-of-fit, we find that the SM has a $\chi^{2}$ per degree of freedom of 0.76, whereas the SMEFT fit yields a significantly improved value of 0.56.222When determining the number of degrees of freedom we have considered each bin as an observable even though this value may be reduced due to correlations among the bins.
The tightest bounds are obtained on the coefficients of two-fermion operators, where 95% probability limits range from 0.25 TeV -2 for $C_{tW}/\Lambda^{2}$ to 5 TeV -2 for $C_{t\varphi}$. These bounds are driven by precise differential cross section measurements in top-quark decay, single top-quark production and associated production processes $\bar{t}H, tt$$\bar{t}Z$ and $t\bar{t}\gamma$. In particular, the studies of rare associated top-quark production processes have evolved from observation to precision measurements. It is illustrative to contrast our result with the first fits of these coefficients at the start of Run 2 in Ref. Buckley et al. (2016), where the data still provided insufficient constraints for a global analysis.
@@
The fit does not yield meaningful bounds on two-quark two-lepton operators, as the measurements of Tab. 3 have insufficient sensitivity for the four degrees of freedom, and therefore these WC have not been considered for this fit. Indeed, even when setting the NP scale as low as 1 TeV the constraints of these operators in the global fit are too weak to lay within the perturbative regime. Dedicated analyses of $t\bar{t}l^{+}l^{-}$ at the LHC with $m_{ll}$ above the $Z$-boson mass window may provide additional constraints on these operators. The analysis of Ref. The CMS collaboration (2023) finds bounds of ${\mathcal{O}}(1~{\mathrm{Te\kern-1.00006ptV}}\,^{-2})$. These bounds, however, strongly rely on terms proportional to $\Lambda^{-4}$. The Run 2 data set cannot constrain these operator coefficients in a global analysis based on a linear parametrisation.
@@ 4.3 Associated production with additional charged leptons
The analysis considers several degrees of freedom that correspond to operators with two charged leptons and two top quarks, listed in the lowermost box of Tab. 2. The LHC experiments can constrain their WC through the study of top-quark production in association with additional charged leptons. Bounds of order one are reported from a SMEFT fit to LHC Run 2 data by the CMS experiment The CMS collaboration (2023).
The projection of HL-LHC bounds on these operator coefficients is based on an extrapolation of the measurements in $pp\rightarrow t\bar{t}Z$ production The ATLAS collaboration (2021). The analysis targets events with three or four isolated charged leptons (electrons and muons). Instead of the events with $m_{ll}\sim m_{Z}$ used in Ref. The ATLAS collaboration (2021), here we focus on off-shell events, where the invariant mass of the di-lepton system lies outside the $Z$-boson mass window. Events are binned in four $m_{ll}$ bins: [100-120] GeV , [120-140] GeV , [140-180] GeV and $m_{ll}>$ 180 GeV  A. Gutierrez Camacho, M. Moreno Llácer and M. Vos (2022). The observed event yields of the Run 2 analysis range from $\sim$ 30 events in the lowest bin to about 10 events for the highest bin. Individual bounds based on Run 2 data with quadratic parametrisations of $\Lambda^{-2}$ and $\Lambda^{-4}$ terms yield results similar to Ref. The CMS collaboration (2023), where the sensitivity is driven by the $\Lambda^{-4}$ terms and most pronounced in the highest mass bin. [^3] Shortly after the release of this manuscript the ATLAS collaboration published a measurement and EFT interpretation of the $t\bar{t}l^{+}l^{-}$ process Aad et al. (2025). Our estimates and conclusions are compatible also to those ATLAS results.
@@ 4.4 Projections
Marginalised bounds remain significantly less precise than individual bounds, even after the complete HL-LHC program, due to “blind directions”, i.e. unresolved correlations between the large number of degrees of freedom. This is confirmed in fits to the top-quark sector by other groups Brivio et al. (2020); Hartland et al. (2019) and even in global Higgs/EW/top fits Celada et al. (2024); Ethier et al. (2021); Ellis et al. (2021). The blind directions have a less pronounced effect in quadratic fits, where dimension-six-squared terms proportional to $\Lambda^{-4}$ are included Ethier et al. (2021), as strictly positive contributions cannot cancel each other.
@@ 5 Prospects for $e^{+}e^{-}$ colliders
Machine P($^{+}, ee^{-}$) Energy Luminosity Reference
ILC $(\pm 30\%,\,\mp 80\%)$ 250 GeV 2 ab-1 A. Aryshev et al. () (editors)
@@ 3 TeV 5 ab-1
FCC-eeee Unpolarised Z-pole 205 ab-1 Benedikt et al. (2025)
240 GeV 10.8 ab-1
350 GeV 0.42 ab-1
365 GeV 2.70 ab-1
@@ 5.4 Projections
To assess the potential of electron-positron colliders, the fit is repeated by adding the projected measurements in $^{+}e^{-}\rightarrow b\bar{b}, e+e$$^{+}e^{-}\rightarrow t\bar{t}$ and $e^{+}e^{-}\rightarrow t\bar{t}H$ production. These processes are accessible in successive energy stages of the electron-positron collider. The impact of each process and energy stage is made clear for the example of the ILC in Fig. 7.
@@
The projections for the top-quark Yukawa, translating $\delta y_{t}=-\frac{v^{2}}{\Lambda^{2}}C_{t\varphi}$, are also presented in Tab. 6. Numbers for lepton colliders are based on an extrapolation in Ref. de Blas et al. (2022a) of detailed studies in Refs. Price et al. (2015); Abramowicz et al. (2019). We emphasise the effect of increasing the ILC from 500 GeV to 550 GeV at the same integrated luminosity of 4 ab-1. The increase in centre-of-mass energy produces a 3.5 factor increase in the cross section which translates into almost a factor of 2 improvement on the individual limit of the Yukawa coupling determination using exclusively the ILC data. Once the HL-LHC data is also considered, the improvement becomes slightly milder, around a 1.4 factor, as shown in Tab. 6.
Higher-energy operation can improve the limit further, as demonstrated by the ILC1000 scenario, with 8 ab-1.
@@
$\delta y_{t}$ Global 16% 5.6% 4.1% 2.9% 1.5% 3.0%
Individual 13% 4.1% 3.4% 2.5% 1.5% 2.7%
Uncombined with 6.3% 3.3% 1.6% 3.6%
HL-LHC Indv.
Table 6: Uncertainties for the top-quark Yukawa coupling at 68% probability for different scenarios, in percent. The ILC500, ILC550 and CLIC scenarios also include the HL-LHC, in the top rows. The ILC1000 scenario includes also ILC500 and HL-LHC, for the top rows. Numbers for lepton colliders are based on an extrapolation in Ref. de Blas et al. (2022a) of detailed studies in Refs. Price et al. (2015); Abramowicz et al. (2019).
@@ 6 Lepton colliders in the 3–30 TeV  range
In this section, we study the impact on the top-quark sector of the SMEFT of lepton colliders operated at very high energy, with $\sqrt{s}$ in the range from several to several tens of TeV. This regime could be unlocked by a muon collider, that presents unique benefits (and technological challenges) in comparison with classical electron-positron machines. The Snowmass process provides several compelling reasons to pursue the construction of a Muon Collider Narain et al. (2022), and the P5 Asai et al. (2023) report has endorsed R&\&D efforts. Unlike circular electron-positron colliders, muon colliders can reach very high centre-of-mass energies without significant energy losses due to synchrotron radiation.
@@ 7 Conclusion
We also provide projections for the high-luminosity phase of the Large Hadron Collider and future colliders. HL-LHC projections are obtained adopting the “S2” scenario that is also used for Higgs measurements Cepeda et al. (2019). In this scenario, experimental uncertainties are strongly reduced $(\propto 1/\sqrt{L_{\text{int}}})$, while theory and modelling uncertainties see a more modest improvement (by a factor 1/2). The bounds improve by a factor 2-3 for most of the operator coefficients. In the projections, we extend previous studies Durieux et al. (2019) by considering also seven operators with two charged leptons and two top quarks, enlarging the total number of Wilson coefficients to 29. In principle, these can be constrained by measurements of $t\bar{t}l^{+}l^{-}$ production with $m_{ll}>m_{Z}$. However, we find that no meaningful global bounds are obtained with a parametrisation based on terms proportional to $\Lambda^{-2}$ .
A lepton collider operated above the $t\bar{t}$} production threshold has excellent potential to constrain the coefficients of bottom- and top-quark operators. The projections used in this paper are based on full-simulation study of optimal observables in Ref. Durieux et al. (2018), with up-to-date operating scenarios for electron-positron Higgs/top/EW factories. Individual bounds on the coefficients $C_{tZ}$ and $C_{tW}$ of the dipole operators reach and on $C_{\varphi Q}^{(1)}$ and $C_{\varphi Q}^{(3)}$ reach 10-2 TeV -2 with a single run with an integrated luminosity of 4 ab-1 at $\sqrt{s}=$ 500 GeV . The bounds on the two-lepton-two-heavy-quark operators even reach 10-3 TeV -2. The global bounds are limited by degeneracies between the two-fermion and four-fermion operators, that can be lifted to a good extent by including measurements at two different centre-of-mass energies. Constraints on the operators considered in our analysis from a possible future lepton collider have also been recently derived in Refs. Maura et al. (2025); Hoeve et al. (2025).
@@ Acknowledgments
We thank J. Tian and M. Peskin for their comments and suggestions on the first version of the manuscript. The authors acknowledge the work of Master students Abel Gutierrez Camacho, Belén Durán González and Pablo Copete Garrido in the development of this analysis. F.C.G is supported by the Presidential Society of STEM Postdoctoral Fellowship at Case Western Reserve University and by the Ministerio de Ciencia, Innovación y Universidades, Spain, through a Beatriz Galindo Junior grant BG23/00061. The work of VM has been supported by the Italian Ministry of Research (MUR) under the grant PRIN20172LNEEZ, by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 949451) and a Royal Society University Research Fellowship through grant URF/R1/201553. The work of MML is supported by the Science and Technology Facilities Council [grant number ST/X005941/1]. MMLL acknowledges the support received from the Spanish Ramón y Cajal programme (RYC2019-028510-I). Her work is also supported by the projects ASFAE/2022/010 and CIPROM/2022/70 (Generalitat Valenciana), and PID2021-124912NB-I00 (Spanish Ministry MICIN). The work of MV is supported by the Spanish ministry for science under grant number PID2021-122134NB-C21, by the Generalitat Valenciana under PROMETEO grant CIPROM/2021-073, and by CSIC under grant ILINKB20065. The group members of IFIC in Valencia received support from the Severo Ochoa excellence programme.